# SAT: Dynamic Spatial Aptitude Training for Multimodal Language Models

### v2

@@ A.3 Dataset Details
Normalizing the camera coordinates In ProcTHOR Deitke et al. (2022), in the camera view, the $y$ coordinate is the height coordinate, which means the $y$ increases pointing upwards (e.g., the ceiling has a greater $y$ than the floor). Hence, from the bird’s eye view, the coordinates are $x$ and $z$. The rotation of the camera is such that it is always parallel to the $x$-$z$ plane. Hence the rotation is described as angle clockwise around the $y$-axis with the camera pointing to the positive z-axis as a 0-degree rotation.
@@
Egocentric Movement. We first choose an image frame. Next, we first choose to rotate left or right from angles 20, 30, 40, 50, 60 chosen randomly. We use the controller.step(action=’RotateRight’, degrees=angle) function in the AI2THOR Kolve et al. (2017) platform. Next, we move forward with probability $0.5$ by a random distance from 20 to 40 centimeters (controller.step(action=’MoveAhead’, moveMagnitude=dist). We capture the next frame from this final position. We also formulate camera movement questions. Note that camera movement and rotation are separate since a camera can be moving left while rotating right. For camera movement, we follow a translate the camera left/right in a random direction by a random distance. Specifically, we first rotate by an angle between 45 to 90 degrees to the left/right, then move forward by a random distance and then rotate right/left (right if previously rotated left in the first step and vice versa) by an angle between 90 and 135 degrees. Sometimes, we do not move or rotate.
Object Movement. We first choose an object visible in the scene with at least a certain bounding box area to make sure the object is not too small or non-salient. Next we decide to move the object by a random distance sampled from 0.25 to 0.5 meters in a random direction if possible. We use the PlaceObjectAtPoint function in the AI2THOR Kolve et al. (2017) platform. Sometimes, we do not move any object. Sometimes the camera as well as an object moves. The agent needs to disentangle the two to answer questions correctly.
Allocentric Perspective. We choose a point on the 2D image that is not too close to the margsns (in between $0.2$-$0.8$ normalized width and height of the image). We use GetCoordinateFromRaycast action in AI2Thor Kolve et al. (2017) to get the 3D location of the point. We try a few points since some points may not possible to navigate to and pick one possible point. We randomly turn left or right by 90 degrees. Next, we check whether an object is left/right based on the method described in Section A.3.
Goal Aiming. We compute the angle between a randomly chosen object with the camera assuming looking straight is 0 degrees. If the angle of an object is $-\alpha$, we say we have to “turn left by $\alpha$ degrees” and “right by $\alpha$ degrees” otherwise. When $|\alpha-0|\leq\epsilon$, we say we have to look “roughly straight”. We define $\epsilon=10$. Note it is very hard for a human or machine to judge the exact degrees to turn from one single image, hence, we give multiple choices where the difference in choices are between left/right and hence, the machine really only has to decide between them and not the exact angles to turn. We use the following equation to calculate the angle of an object with the camera. First we normalize the object coordinates to the camera coordinates, $(x_{0},y_{0})$, and then calculate the angle, $\alpha$.
@@
We use DepthAnything (Yang et al., 2024b) to estimate depth of the scene and bounding box annotations in GQA (Hudson & Manning, 2019a) and VG (Krishna et al., 2017) to formulate static spatial QAs. To make spatial QAs, we first choose three objects in the scene based on which we make questions as defined in Section 3.1. We use the following heuristics to first choose the three objects to reduce noise:
@@
For VSR (Liu et al., 2023), we have captions like “truck in front of airplane.” We re-formulate them to into quetsion-answer pairs like: “Is truck in front of airplane? Answer yes or no” to keep all evaluation formats consistent. This ensures performance differences are due to the difficulties of spatial relationships in the datasets and not the question-answering format.
@@
For VSI-Bench (Yang et al., 2024a), we only use the MC split and not the numerical splits. We believe the numerical splits requiring estimation of exact distances and dimensions from RGB images alone without camera and depth information is difficult and out of scope for SAT since SAT focuses on relationships in images and dynamic scenes and not exact depth and distances.
@@
Table 8: Datasource for LLaVA-VIdeo on long video VSI-Bench
@@
LLaVA-Video 43.9 42.0 33.5 32.4 38.0
+ SAT Static 43.5 43.4 34.5 30.3 37.9
+ SAT Dynamic 47.9 39.6 38.7 40.6 41.7
@@
We frame each questions as a binary choice following existing benchmark standards (Fu et al., 2024b; Tong et al., 2024), and to prevent the MLM from choosing shortcuts by eliminating obvious wrong choices in multiple choices (Cai et al., 2024).
Following LLaVA Liu et al. (2024) convention, we use <image> tokens to represent images. This is the exact prefix we use.
@@
Errors in relationship judgment compounded with lack of reasoning causes MLMs to perform worse than random on dynamic tasks For certain splits such as perspective, or overall performance on SAT real, some accuracies are below random chance (random is $50\%$). In dynamic examples, the spatial position of an object often switches when a new perspective is taken—e.g., a cup might be in the left side of the image, but switches to the right side after taking a new perspective. Weaker MLMs have bad spatial perception in general. However, even for MLMs with strong static relationship performance, they often answer what is currently visible (i.e., cup is on the left) rather than what actually happens after the new perspective (i.e., cup on the right). Examples are shown in Figure 5. For example, let’s look at the performance of a model which has strong static spatial performance - Gemini-1.5-pro. If we look at the second row example, the surfer is currently on the left of the X marked point and hence the model answers left instead of reasoning about how it might switch when viewing from that perspective. Similarly, in row 3, and in row 7 where the door is currently visible on the left.
Adding in dynamic helps LLaVA-Video on dynamic and video spatial reasoning As shown in Tables 7 and 8, adding in dynamic data helps LLaVA-Video on video spatial reasoning (Table 8) and on dynamic SAT Real in Table 7, while maintaining performance on CVBench and BLINK. While scaling works better with mixing in dynamic as shown in Figure 4 in the main paper, we also analyze the effect of adding dynamic data while maintaining the same data size as static in Table 9. We see that for both models, adding in dynamic helps significantly on SAT Real. For LLaVA1.5 it helps overall. For LLaVA-Video, it helps on video and SAT real while maintaining performance on BLINK and CVBench.
Quality of data is important. Hence, simulations are an effective source. We wish to answer if the gains trivially come from just simply learning the formats. Hence, we run a noise baseline where we flip the answers $20\%$ of the time. We see performance degrades significantly. We see a reduction of $10\%$ on CVB and $18\%$ on BLINK when training with just $20\%$ noise. This means MLMs are not trivially learning the format of questions or biases and rather the visual input and answer quality is important.

### v3

@@ A.3 Dataset Details
Normalizing the camera coordinates In ProcTHOR deitke2022️, in the camera view, the $y$ coordinate is the height coordinate, which means the $y$ increases pointing upwards (e.g., the ceiling has a greater $y$ than the floor). Hence, from the bird’s eye view, the coordinates are $x$ and $z$. The rotation of the camera is such that it is always parallel to the $x$-$z$ plane. Hence the rotation is described as angle clockwise around the $y$-axis with the camera pointing to the positive z-axis as a 0-degree rotation.
@@
Egocentric Movement. We first choose an image frame. Next, we first choose to rotate left or right from angles 20, 30, 40, 50, 60 chosen randomly. We use the controller.step(action=’RotateRight’, degrees=angle) function in the AI2THOR Kolve2017AI2THORAn platform. Next, we move forward with probability $0.5$ by a random distance from 20 to 40 centimeters (controller.step(action=’MoveAhead’, moveMagnitude=dist). We capture the next frame from this final position. We also formulate camera movement questions. Note that camera movement and rotation are separate since a camera can be moving left while rotating right. For camera movement, we follow a translate the camera left/right in a random direction by a random distance. Specifically, we first rotate by an angle between 45 to 90 degrees to the left/right, then move forward by a random distance and then rotate right/left (right if previously rotated left in the first step and vice versa) by an angle between 90 and 135 degrees. Sometimes, we do not move or rotate.
Object Movement. We first choose an object visible in the scene with at least a certain bounding box area to make sure the object is not too small or non-salient. Next we decide to move the object by a random distance sampled from 0.25 to 0.5 meters in a random direction if possible. We use the PlaceObjectAtPoint function in the AI2THOR Kolve2017AI2THORAn platform. Sometimes, we do not move any object. Sometimes the camera as well as an object moves. The agent needs to disentangle the two to answer questions correctly.
Allocentric Perspective. We choose a point on the 2D image that is not too close to the margsns (in between 0.20.2-$0.8$ normalized width and height of the image). We use GetCoordinateFromRaycast action in AI2Thor Kolve2017AI2THORAn to get the 3D location of the point. We try a few points since some points may not possible to navigate to and pick one possible point. We randomly turn left or right by 90 degrees. Next, we check whether an object is left/right based on the method described in Section A.3.
Goal Aiming. We compute the angle between a randomly chosen object with the camera assuming looking straight is 0 degrees. If the angle of an object is $\alpha$, we say we have to “turn left by $\alpha$ degrees” and “right by $\alpha$ degrees” otherwise. When $|\alpha-0|\leq\epsilon$, we say we have to look “roughly straight”. We define $\epsilon=10$. Note it is very hard for a human or machine to judge the exact degrees to turn from one single image, hence, we give multiple choices where the difference in choices are between left/right and hence, the machine really only has to decide between them and not the exact angles to turn. We use the following equation to calculate the angle of an object with the camera. First we normalize the object coordinates to the camera coordinates, $(x_{0},y_{0})$, and then calculate the angle, $\alpha$.
@@
We use DepthAnything (depthanything) to estimate depth of the scene and bounding box annotations in GQA (hudson2019gqanewdatasetrealworld) and VG (krishna2017visual) to formulate static spatial QAs. To make spatial QAs, we first choose three objects in the scene based on which we make questions as defined in Section 3.1. We use the following heuristics to first choose the three objects to reduce noise:
@@
For VSR (liu2023visualspatialreasoning), we have captions like “truck in front of airplane.” We re-formulate them to into quetsion-answer pairs like: “Is truck in front of airplane? Answer yes or no” to keep all evaluation formats consistent. This ensures performance differences are due to the difficulties of spatial relationships in the datasets and not the question-answering format.
@@
For VSI-Bench (yang2024thinking), we only use the MC split and not the numerical splits. We believe the numerical splits requiring estimation of exact distances and dimensions from RGB images alone without camera and depth information is difficult and out of scope for SAT since SAT focuses on relationships in images and dynamic scenes and not exact depth and distances.
@@
Table 9: Datasource for LLaVA-VIdeo on long video VSI-Bench. Row b and c are trained purely on simulations, while row d is simulation and orignal pre-training data (LLaVA-OV data) mixed in during training to prevent forgetting.
@@
a. LLaVA-Video 43.9 42.0 33.5 32.4 38.0
b. + SAT Static 43.5 43.4 34.5 30.3 37.9
c. + SAT Static + Dyn 47.7 42.0 35.6 36.1 40.4
d. + SAT Dyn + LLaVA-OV 47.9 39.6 38.7 40.6 41.7
@@
Table 11: Most frequent correct and incorrect words
Rank Correct Words Incorrect Words Incorrect Words (excluding count)
1 stationery 3 4x4
2 pan 5 iced
3 books 4 wheeler
4 my 4x4 donuts
5 into 12 cookies
6 vase 6 tea
7 straight iced poles
8 magazine 10 office
9 signs 11 game
10 flowerpot 8 dogs
@@
We frame each questions as a binary choice following existing benchmark standards (fu2024blink; tong2024cambrian1fullyopenvisioncentric), and to prevent the MLM from choosing shortcuts by eliminating obvious wrong choices in multiple choices (cai2024temporalbench).
Following LLaVA liu2024visual convention, we use <image> tokens to represent images. This is the exact prefix we use.
@@
Analyzing the semantic concepts where models fail after SAT tuning. We wish to understand the semantic concepts where SAT tuning doesn’t suffice. Hence, we we analyze the words that most often occur in the failure cases in contrast to the correct cases and show them in Table 11. Unsurprisingly, we see a high fraction of count words with higher numbers than 2 being mostly present. If we exclude the coutn words, we see a higher fraction of outdoor words. This is also unsurprising since our simulator only has indoor scenes. Using world models and more diverse simulators to improve our dataset is an exciting area of future work.
Errors in relationship judgment compounded with lack of reasoning causes MLMs to perform worse than random on dynamic tasks For certain splits such as perspective, or overall performance on SAT real, some accuracies are below random chance (random is $50\%$). In dynamic examples, the spatial position of an object often switches when a new perspective is taken—e.g., a cup might be in the left side of the image, but switches to the right side after taking a new perspective. Weaker MLMs have bad spatial perception in general. However, even for MLMs with strong static relationship performance, they often answer what is currently visible (i.e., cup is on the left) rather than what actually happens after the new perspective (i.e., cup on the right). Examples are shown in Figure 5. For example, let’s look at the performance of a model which has strong static spatial performance - Gemini-1.5-pro. If we look at the second row example, the surfer is currently on the left of the X marked point and hence the model answers left instead of reasoning about how it might switch when viewing from that perspective. Similarly, in row 3, and in row 7 where the door is currently visible on the left. While this analysis is from qualitative analysis, a more thorough analysis is left as future work.
Adding in dynamic helps LLaVA-Video on dynamic and video spatial reasoning As shown in Tables 8 and 9, adding in dynamic data helps LLaVA-Video on video spatial reasoning (Table 9) and on dynamic SAT Real in Table 8, while maintaining performance on CVBench and BLINK. While scaling works better with mixing in dynamic as shown in Figure 4 in the main paper, we also analyze the effect of adding dynamic data while maintaining the same data size as static in Table 10. We see that for both models, adding in dynamic helps significantly on SAT Real. For LLaVA1.5 it helps overall. For LLaVA-Video, it helps on video and SAT real while maintaining performance on BLINK and CVBench.
Quality of data is important. Hence, simulations are an effective source. We wish to answer if the gains trivially come from just simply learning the formats. Hence, we run a noise baseline where we flip the answers $20\%$ of the time. We see performance degrades significantly. We see a reduction of $10\%$ on CVB and $18\%$ on BLINK when training with just $20\%$ noise. This means MLMs are not trivially learning the format of questions or object/attribute -name-based biases. Rather, the spatial answer quality is important.